solve the equation 2x^2-3x+1=0
and present solutions in solutions set notation

Answer by checkley77(9423)

2x^2-3x+1=0
(2x-1)(x-1)=0
2x-1=0
2x=1
x=1/2 ans.
x-1=0
x=1 ans.